OpenStudy (anonymous):

bob plays a game at the school fair. he starts with 0 points. he gets 25 throws. he wins 12 points for hitting the target and loses 8 points for each miss. bob ends with a score of 0. how many hits and misses does bob have?

OpenStudy (anonymous):

show work

OpenStudy (anonymous):

how much does he lose per miss?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

So... Hits = h Misses = m 12h - 8m = 0, or 12h = 8m h + m = 25

OpenStudy (anonymous):

then solve the equations

OpenStudy (anonymous):

simple

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

12h - 8m = 0, which means that 12h = 8m.

OpenStudy (anonymous):

h + m = 25

OpenStudy (anonymous):

So m = 25 - h

OpenStudy (anonymous):

Replace m with (25 - h) in the equation 12h = 8m

OpenStudy (anonymous):

then solve for 12h = 8(25-h)

OpenStudy (anonymous):

thats 10 hits, but how about misses?

OpenStudy (anonymous):

There is 25 total... There is 10 misses... Simple Subtraction.

OpenStudy (anonymous):

oh... lol .. huge fail for me thanks

OpenStudy (anonymous):

np